# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v2

@@ 3.5. Walkthrough of Revising Step
3.5.1. Frontend
In the Revising-Support Area (Figure 2d), the user works with the tool to revise the blog post section on which they are currently working (DG3). At the top of this area, the user pastes the text that they want to modify in the “Writing to Modify” text field. This text may be anything, including text from the blog post draft, the paper, or a previous model generation. The user then adjusts the desired length of the modified text if need be; the options are the same as in the Drafting-Support Area (Section 3.4). Also as in the Drafting-Support Area, the user may adjust the first toggle, which is on by default to indicate that the blog post sections above the one being edited are being provided as context to the LLM. The user then notes the second toggle, which is on by default to indicate that the LLM will utilize the currently selected paragraphs and bullet points to generate the modified text. At the bottom of the Revising-Support Area, there are five preset modification buttons to 1) expand the text, 2) condense the text, 3) rewrite the text in simpler terms to be more understandable to a layperson, 4) rewrite the text in a less dramatic tone to align better with the unadorned language of scientific writing, and 5) rewrite the text in a more dramatic tone to better capture readers’ attention. There is also a “custom modification” button. If the user wants to use the “custom modification” button, they enter their custom instructions for the text modification in the text field above. If the user wants to use a different modification button, they may still provide custom instructions for the LLM to take into consideration. The user copies and pastes the modified text into the highlighted blog post section and edits it as they see fit. The user may peruse a history of their LLM instructions and text modifications for the highlighted section.
3.5.2. Backend
As shown in Figure 1d, the LLM (GPT-4 in lab study; GPT-4-0125-preview in deployment) modifies text for the highlighted blog post section with the user’s selected text snippet and their adjusted instructions for the modification (described in Section 3.5.1), including the selected text modification. The desired length “auto” is unique to each modifying action. The prompts for revising text may be found in Appendix B.4, and example outputs are in Appendix F.3.
@@ 4.1. Lab Study
We conducted a lab study to determine if and how Papers-to-Posts may support researchers in writing blog posts about their research papers.
@@ 4.1.1. Hypotheses
  • H4: Papers-to-Posts leads to less time spent generating a blog post that the writer would be comfortable posting publicly, in comparison to the baseline.
@@ 4.1.2. Participants
We recruited 20 participants (M: 14, W: 6) via academic social networks and institutional mailing lists and compensated them with \$83.33 USD for 2.5 hours of their time. Participants were predominantly early career researchers (PhD student: 17, master’s student: 1, postdoc: 1, industry researcher: 1), who are often lead authors tasked with writing blog posts. All participants conducted research in computer science. Eight participants noted being in a subarea related to human-computer interaction. Participants had a broad range of experience in terms of prior LLM use (¿10 sittings: 14, 6-10 sittings: 3, 2-5 sittings: 2, ¡2 sittings: 1) and number of research-paper blog posts authored (0 posts: 14, 1-2 posts: 5, 3-10 posts: 1). In order to simulate a naturalistic setting, participants were required to have authored at least 2 research papers for use in the study (2 papers: 1, 3-10 papers: 14, ¿10 papers: 5) and to be more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing blog posts for their research papers.
@@ 4.1.3. Study Conditions
The treatment condition involved interacting with Papers-to-Posts  which is described in Section 3. The tool in the baseline condition was designed to be a simplified version of Papers-to-Posts without affordances developed to address the design goals (Figure 4). The baseline provided input paper text, sourced from the PDF and divided into paragraphs and collapsible sections as in Papers-to-Posts. The baseline also provided the ability to prompt GPT-4 with any instructions and copy its output. A history of generations and their associated instructions was retained for the user’s perusal. Lastly, the baseline had an editor for writing the blog post, which was pre-populated with an LLM draft generated by GPT-4-32k using the entirety of the input paper. [^4] An example of a baseline initial draft is in Appendix E. GPT-4-32k was used rather than GPT-4 because the context window for GPT-4 was insufficient for longer research papers. The prompt was designed to mirror the prompt for Papers-to-Posts’s initial draft sections (Appendix B.1), so the resulting draft also had introduction, methods, results, and conclusion sections. However, the prompt did not contain section-specific instructions.
@@
In each condition, the tool was set up for one of the participant’s research papers submitted in the recruitment survey. The paper was different for each condition. Each paper was required to have been published online between 2018 and the study period in 2023, with the participant as an author. Only one paper was published in 2018, and the rest were published during or after 2020. There was no significant difference in publication dates between conditions (Paired-Samples t-Test, t(19)=0.90, p=n.s.). The average paper length across conditions differed by less than 10% (baseline: 8452 words, treatment: 7742 words).
@@ 4.1.4. Study Procedure
Participants completed a participation agreement before the study. We recorded and transcribed each session in Google Hangouts. [^6] The lab study session script may be found in the supplementary materials. To start, the session coordinator introduced the task and presented a video tutorial on the assigned tool. [^7] See the tutorial scripts in the supplementary materials. Participants then had two minutes to interact with and ask questions about the tool in the context of a sample paper that was not their own. Next, participants opened a tool link initialized with one of their research papers. The coordinator asked them to assume that their audience was people in their broad area (i.e., computer science), which was the most common audience for whom potential study recruits wanted to write blog posts (52/60). The coordinator instructed participants to work as if they were not there, but to let them know if any questions about the tool arose. They also instructed participants not to use any external tools or information. We made an exception twice when participants wanted to view their paper’s PDF due to issues with its parsing in the tool. Lastly, the coordinator asked participants to let them know if they reached a point at which they would feel comfortable publicly publishing the blog post in its current state.
@@
Each participant had up to one hour to work on the blog post. At the 20-minute and 40-minute marks, they were reminded to let the coordinator know if they reached the point of comfort with publication. If and when the participant reached that point, the coordinator asked them to save their draft and continue revising the blog post. After the first four study sessions, we noticed that participants were sometimes done early and did not desire to revise further. Thus, we added an additional rule that the participant would let the coordinator know if they reached a point at which they were proud of the blog post and saw no reason to revise further. If and when the participant reached that point, or if an hour had passed, they submitted their final blog post and moved on to a survey regarding their experience with the tool. The survey [^8] The survey may be found in the supplementary materials. consisted of Likert-type questions regarding perceived efficiency, cognitive load, ease of using the tool in relation to the design goals, satisfaction with the tool, satisfaction with the final blog post, perceived control, and perceived ownership. [^9] See Appendix A for a discussion of the exploratory results regarding participants’ perceived control and ownership. There were also short answer questions regarding the difficulties and benefits of using the tool in relation to the design goals. [^10] If it was the participant’s second session, the survey ended with a few questions directly comparing the two conditions. These were not for analysis but only to help us get a quick sense of the results. At the session’s end, the coordinator told the participant to close the tool and not to access it before the next session if they had one. Each time the participant prompted the LLM or submitted the blog post, a log of the action along with the state of the blog post was saved for analysis.
@@ 4.2.1. Research Questions
  • RQ1: How is potential users’ editing power impacted by Papers-to-Posts versus the baseline tool in a real-world setting?
  • RQ2: How do potential users perceive the outcome (blog post satisfaction (a)) and experience (tool satisfaction (b), cognitive load (c)) of interacting with Papers-to-Posts versus the baseline tool in a real-world setting?
@@ 4.2.2. Participants
We recruited 26 participants (M: 17, W: 8, undisclosed: 1) through academic social networks and institutional mailing lists. They were compensated with \$50 over PayPal for each blog post that they wrote, with a maximum of two blog posts. Participants were predominantly early career researchers (undergraduate student: 1, master’s student: 4, PhD student: 13, postdoc: 2, professor: 2, industry researcher: 4). All participants conducted research in an area of computer science, except for one who studied cognitive science. The majority of the participants had interacted with an LLM several times (¿10 sittings: 21, 6-10 sittings: 4, 2-5 sittings: 1). The participants had varied experience in writing blog posts (0 posts: 16, 1-2 posts: 9, 3-10 posts: 1). They were required to have authored at least one research paper (1-2 papers: 6, 3-10 papers: 11, ¿10 research papers: 9) and to be more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing research-paper blog posts. We also required participants to be willing to share their final blog posts with others if they were able to reach the point at which they would feel comfortable doing so. Lastly, the papers that the participants want to write about had to be compatible with our tools’ PDF parser.
@@ 4.2.3. Study Conditions
While the treatment and baseline conditions were largely the same as in the lab study, several minor usability issues identified in the lab study were addressed before proceeding with the deployment study. For example, there were better instructions and less intrusive highlight colors (Appendix 28). We also replaced GPT-4 and GPT-4-32k with the newer GPT-4-0125-preview model, which had a larger context window and became available after the lab study. This enabled Papers-to-Posts to support longer inputs. In the deployment study, participants could freely close and reopen their assigned tools in their browsers to work on their drafts when convenient. Participants could save their progress across sessions with a newly added save button. We note that, due to the manner in which text was logged in the deployment study, the tool in both conditions had a lag if participants pasted a chunk of text more then a couple paragraphs long. Lastly, the treatment tool now automatically parsed the input paper.
@@ 4.2.4. Study Procedure
The task for the deployment study was to share a blog post with others, which can require substantially higher effort than voicing comfort in publicly posting the blog post, as in the lab study. Considering this difference for participant recruitment, we opted for a between-subjects design, so each participant only had to write one blog post. Still, we allowed participants who wished to participate twice (N=11/26 participants) to use both tools in randomized, counter-balanced order. Their selected papers were assigned randomly to each condition. When analyzing these participants’ results, we treated them as unpaired data. Eighteen participants completed the baseline, and 19 completed the treatment. Participants had up to six days to complete a paper blog post using the assigned tool on their own time. We provided a link to the tool, a video tutorial on the tool, and a link to the paper about which to write. We noted that participants should aim to share the final blog post with others, but if they felt that they could not reach that point, they would still be compensated. Participants were required to spend at least 20 minutes using the tool. We recorded participants’ interactions, including prompting and text edits. In addition, we also took snapshots of the blog post whenever they (re)opened the systems in their browsers and every 60 seconds during active usage.
@@
Once participants were satisfied with their draft, they shared it. We provided instructions on how to create blog posts on the Medium platform, [^12] https://medium.com/ but they were allowed to use other publishing platforms if they preferred. They then emailed us the blog post link and a screenshot or link showcasing evidence that they shared the blog post with others (e.g., tweet, email to co-authors, Slack channel message, LinkedIn post). Only 2 of the 26 participants (one under both conditions and one under the treatment) could not reach the point at which they felt comfortable sharing the blog post with others in the allotted time. [^13] We note that some participants indicated that the blog post was part of an experiment when sharing with others, which may have increased their willingness to share it; regardless, they were willing to share a post at the end of the study. Participants then completed the same survey from the lab study regarding their experience (see Section 4.1.4). [^14] We note that four participants who we realized had not completed 20 minutes of interaction were asked to do so after completing the survey. The only difference was the removal of a few questions irrelevant to this study151515a question regarding how much additional time would be needed to feel comfortable posting the blog post publicly and questions regarding a direct comparison between the two tools. Lastly, the first author engaged willing participants (15) in a 15-minute semi-structured interview to discuss their experience and survey responses.

### v3

@@ 3.6. Walkthrough of Revising Step
In the Revising-Support Area (Figure 3d), the user works with the tool to revise the blog post section on which they are currently working, addressing DG3. At the top of this area, the user pastes the text that they want to modify in the text field. This text may be anything, including text from the blog post draft, the paper, or a model generation. The user then adjusts the desired length of the modified text if need be, as in the Drafting-Support Area (Section 3.5). [^8] The desired length “auto” is unique to each modifying action. The user leaves the second toggle on, indicating that the LLM will utilize the currently selected paragraphs and bullet points to generate the modified text. The user considers the five preset modification options: 1) expand the text, 2) condense the text, 3) rewrite the text in simpler terms to be more understandable to a layperson, 4) rewrite the text in a less dramatic tone to align better with the unadorned language of scientific writing, and 5) rewrite the text in a more dramatic tone to better capture readers’ attention. The preset modification buttons called prompts based on our synthesized guidelines for academic blog posts (see 3.1), which suggested avoiding jargon (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), sensationalism (noted by a communications expert), and wordiness (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]). There is also a “custom modification” button that will follow specified custom instructions. [^9] If the user utilizes a preset modification button, they may still provide custom instructions for the LLM to take into consideration. The user copies and pastes the modified text into the highlighted blog post section and edits it as needed. If the user wants to remember their previous text modifications for this section, they may peruse a history of their instructions and modifications. With the user’s instructions, the LLM modifies the user’s provided text snippet. [^10] The prompts for revising text may be found in Appendix B.4, and example outputs are in Appendix F.3.
@@ 4.1. Lab Study
We conducted a within-subjects lab study to determine if and how Papers-to-Posts may support researchers in writing blog posts about their research papers.
@@ 4.1.1. Hypotheses
  • H4: Papers-to-Posts leads to less time spent to generate a blog post that the writer would be comfortable posting publicly, in comparison to the baseline.
@@ 4.1.2. Participants
We recruited 20 participants (M: 14, W: 6) via academic social networks and institutional mailing lists and compensated them with \$83.33 USD for 2.5 hours of their time. Participants were predominantly early career researchers (PhD student: 17, master’s student: 1, postdoc: 1, industry researcher: 1), who are often lead authors tasked with writing blog posts. All participants conducted research in computer science. Participants had a broad range of experience in terms of LLM use (¿10 sittings: 14, 6-10 sittings: 3, 2-5 sittings: 2, ¡2 sittings: 1). Most had not authored a research-paper blog post before (0 posts: 14, 1-2 posts: 5, 3-10 posts: 1). In order to simulate a naturalistic setting, participants were required to have authored at least 2 research papers for use in the study (2 papers: 1, 3-10 papers: 14, ¿10 papers: 5) and to be more interested than not in writing blog posts for their research papers.
@@ 4.1.3. Study Conditions
In the treatment condition, participants interacted with Papers-to-Posts  described in Section 3.111111The system figures above show how the tool looked for the deployment study, after minor usability updates (see Section 4.2.3. Appendix 28 shows how the tool looked for the lab study. The tool in the baseline condition was designed to be a simplified version of Papers-to-Posts without affordances developed to address the design goals (Figure 4). The baseline provided the input paper’s text, sourced from the PDF and divided into paragraphs and collapsible sections as in Papers-to-Posts. The baseline also provided the ability to prompt GPT-4 with any instructions and copy its output. A history of generations and their associated instructions was retained for the user’s perusal. Lastly, the baseline had an editor for writing the blog post, which was pre-populated with an LLM draft generated by GPT-4-32k using the entirety of the input paper. [^12] An example of a baseline initial draft is in Appendix E. GPT-4-32k was used rather than GPT-4 because the context window for GPT-4 was insufficient for longer research papers. The prompt was designed to mirror the prompt for Papers-to-Posts’ initial draft sections (Appendix B.1), so the resulting draft also had an introduction, methods, results, and conclusion section. However, the prompt did not contain section-specific instructions.
@@
The image may be divided into three areas: one on the left, one on the top right, and one on the bottom right. The left area is titled ”Blogpost.” There is a white text box filling the entire area. It contains text with three sections (introduction, methods, and results) visible. The top right area is titled ”Paper” and shows two section header with a gray background. The text of the sections has a white background. The bottom right area is yellow and titled ”Generate Writing.” It contains a text box labeled ”Instructions*” with placeholder text that says ”Enter any instructions for the AI tool.” Below the text box, there is a button labeled ”Generate.”
In each condition, the tool was set up for one of the participant’s two research papers submitted in the recruitment survey. Each paper was required to have been published online between 2018 and the study period in 2023, with the participant as an author. Only one paper was published in 2018, and the rest were published during or after 2020. There was no significant difference in publication dates between conditions (Paired-Samples t-Test, t(19)=0.90, p=n.s.). The average input paper length across conditions differed by less than 10% (baseline: 8452 words, treatment: 7742 words).
@@ 4.1.4. Study Procedure
Participants completed a participation agreement before the study. We recorded and transcribed each session in Google Meet. [^14] The lab study session script may be found in the supplementary materials. To start, the session coordinator introduced the task and presented a video tutorial on the assigned tool. [^15] The tutorial scripts are in the supplementary materials. Participants then had two minutes to interact with and ask questions about the tool in the context of a sample paper that was not their own. Next, participants opened a tool link initialized with one of their research papers. The coordinator asked them to assume that their audience was people in their broad area (i.e., computer science), which was the most common audience for whom potential study recruits wanted to write blog posts (52/60). The coordinator instructed participants to work as if they were not there, but to let them know if any questions about the tool arose. They also instructed participants not to use any external tools or information. [^16] We made an exception twice when participants wanted to view their paper’s PDF due to issues with its parsing in the tool. Lastly, the coordinator asked participants to let them know if they reached a point at which they would feel comfortable publicly publishing the blog post in its current state.
@@
Each participant had up to one hour to work on the blog post. At the 20-minute and 40-minute marks, they were reminded to let the coordinator know if they reached the point of comfort with publication. If the participant reached that point, the coordinator asked them to save their draft and continue revising the blog post. After the first four study sessions, we noticed that participants were sometimes done early and did not desire to revise further. Thus, we added an additional rule that the participant would let the coordinator know if they reached a point at which they were proud of the blog post and saw no reason to revise further. If the participant reached that point, or if an hour had passed, they submitted their final blog post and moved on to a survey regarding their experience with the tool. The survey consisted of Likert-type questions regarding perceived efficiency, cognitive load, ease of using the tool in relation to the design goals, satisfaction with the tool, satisfaction with the final blog post, perceived control, and perceived ownership. [^17] See Appendix A for a discussion of the exploratory results regarding participants’ perceived control and ownership. There were also short answer questions regarding the difficulties and benefits of using the tool in relation to the design goals. [^18] The survey may be found in the supplementary materials. At the session’s end, the coordinator told the participant to close the tool and not to access it before the next session if they had one. Each time the participant prompted the LLM or submitted the blog post, a log of the action along with the state of the blog post was saved for analysis.
@@ 4.2.1. Research Questions
  • RQ1: How is participants’ editing power impacted by Papers-to-Posts versus the baseline tool in a real-world setting?
  • RQ2: How do participants perceive the outcome (blog post satisfaction) and experience (tool satisfaction, cognitive load) of interacting with Papers-to-Posts versus the baseline tool in a real-world setting?
@@ 4.2.2. Participants
We recruited 26 participants (M: 17, W: 8, undisclosed: 1) through academic social networks and institutional mailing lists. They were compensated with \$50 over PayPal for each blog post that they wrote, with a maximum of two blog posts. Participants were predominantly early career researchers (undergraduate student: 1, master’s student: 4, PhD student: 13, postdoc: 2, professor: 2, industry researcher: 4). All participants conducted research in an area of computer science, except for one who studied cognitive science. The majority of the participants had interacted with an LLM several times (¿10 sittings: 21, 6-10 sittings: 4, 2-5 sittings: 1). The participants had varied experience in writing blog posts (0 posts: 16, 1-2 posts: 9, 3-10 posts: 1). They were required to have authored at least one research paper (1-2 papers: 6, 3-10 papers: 11, ¿10 research papers: 9) and to be more interested than not in writing research-paper blog posts. We also required participants to be willing to share their final blog posts with others if they were able to reach the point at which they would feel comfortable doing so. Lastly, the papers that the participants wanted to write about had to be compatible with our tools’ PDF parser.
@@ 4.2.3. Study Conditions
While the treatment and baseline conditions were largely the same as in the lab study, several minor usability issues identified in the lab study were addressed before proceeding with the deployment study. For example, there were better instructions and less intrusive highlight colors, as depicted in the system figures above. We also replaced GPT-4 and GPT-4-32k with the newer GPT-4-0125-preview model, which had a larger context window and became available after the lab study. This enabled both conditions to support longer inputs. In the deployment study, participants could freely close and reopen their assigned tools in their browsers to work on their drafts when convenient, and they could save their progress across sessions with a newly added save button. Furthermore, the treatment tool now automatically parsed the input paper. We note that, due to the manner in which text was logged in the deployment study, the tool in both conditions had a lag if participants pasted a chunk of text more then a couple paragraphs long.
@@ 4.2.4. Study Procedure
The task for the deployment study was to share a blog post with others, which can require substantially higher effort than voicing comfort in publicly posting the blog post, as in the lab study. Considering this difference for participant recruitment, we opted for a between-subjects design, so each participant only had to write one blog post. Still, we allowed participants who wished to participate twice (N=11/26 participants) to use both tools in randomized, counter-balanced order. Their selected papers were assigned randomly to each condition. When analyzing these participants’ results, we treated them as unpaired data. Eighteen participants completed the baseline, and 19 completed the treatment. Participants had up to six days to complete a paper blog post using the assigned tool on their own time. We provided a link to the tool, a video tutorial on the tool [^20] The scripts for the video tutorials are in the supplementary materials., and a link to the paper about which to write. We noted that participants should aim to share the final blog post with others, but if they felt that they could not reach that point, they would still be compensated. Participants were required to spend at least 20 minutes using the tool. We recorded participants’ interactions, including prompting and text edits, in logs. In addition, we also took snapshots of the blog post whenever they (re)opened the systems in their browsers and every 60 seconds during active usage.
@@
Once participants were satisfied with their draft, they shared it. We provided instructions on how to create blog posts on the Medium platform, [^21] https://medium.com/ but they were allowed to use other publishing platforms if they preferred. They then emailed us the blog post link and a screenshot or link showcasing evidence that they shared the blog post with others (e.g., tweet, email to co-authors, Slack channel message, LinkedIn post). Only 2 of the 26 participants (one under both conditions and one under the treatment) could not reach the point at which they felt comfortable sharing the blog post with others in the allotted time. [^22] We note that some participants indicated that the blog post was part of an experiment when sharing with others, which may have increased their willingness to share it; regardless, they were willing to share a post at the end of the study. Participants then completed the same survey from the lab study regarding their experience (see Section 4.1.4). [^23] We note that four participants who we realized had not completed 20 minutes of interaction were asked to do so after completing the survey. The only difference was the removal of a few questions irrelevant to this study. Lastly, the first author engaged willing participants (15) in a 15-minute semi-structured interview to discuss their experience and survey responses.
(a) Lab Study
(b) Deployment Study
[…]
The top of the figure shows a legend with the following mapping: dark red = strongly disagree, medium red = disagree, light red = somewhat disagree, gray = neither agree nor disagree, light blue = somewhat agree, medium blue = agree, and dark blue = strongly agree. Below, there are two areas. The left area is labeled ”Lab Study,” and the right area is labeled ”Deployment Study.” In each area, there are four questions labeled ”a” through ”d” for the lab study and ”e” through ”h” for the deployment study. The questions are: 1) ”I am confident that all essential information that wanted to cover is in the blog post.”; 2) ”It was easy to incorporate information from the paper in the blog post.”; 3) ”It was easy to iterate on the blog post writing.”; 4) ”It was easy to produce appropriate text for the blog post.” Below each question, there are two horizontal stacked bars, one labeled ”treatment” and the other ”baseline.” In the lab study area, treatment is shown to perform somewhat better than baseline for each question. For all of the questions, the medians for treatment followed by baseline are agree vs. somewhat agree. In the deployment study area, treatment is shown to perform somewhat better than baseline for the second and third question. It appears to do slightly worse overall for the first question, and which condition did better is unclear for the last question. For each question, the medians for treatment followed by baseline are: somewhat agree vs. somewhat agree, agree vs between neither agree nor disagree and somewhat agree, agree vs somewhat agree, and somewhat agree vs. somewhat agree.
5. Evaluation Studies’ Results
5.1. Design Goals 1 and 2
We analyzed participants’ Likert-type responses to questions related to fulfilling our first two design goals, which focused on helping researchers understand and control what content is included in their research-paper blog posts (Figure 5(b)a-c,e-g).
5.1.1. Lab Study - Exploratory
Based on the median results, using Papers-to-Posts, participants were only slightly more confident that all essential information was in the blog post (treatment: M=6.00, Q1=3.00, Q3=6.25; baseline: M=5.00, Q1=3.00, Q3=6.00). They found it easier to incorporate information from the paper (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.75, Q3=6.00), and they found it easier to iterate on the blog post text (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.00, Q3=6.00) (Figure 5(b)a-c). The interaction logs showed that participants utilized Papers-to-Posts’ affordance designed to support these outcomes, selecting relevant paper bullet points and paragraphs from the LLM-generated source outline to incorporate in the blog post. All participants adjusted the LLM’s pre-selected content to some degree. On average, participants added or removed 28 bullet points and 13 paragraphs.  [^24] Some added bullet points come from automatic selection of bullet points used to generate a new section rather than manual selection. These averages do not include 2 participants due to a logging error. Participants also added their own custom bullet points an average of 1.70 times.
5.1.2. Deployment Study - Exploratory
Looking at the median results, participants did not indicate an improvement in feeling confident that all essential information was covered with Papers-to-Posts (M=5.00, Q1=4.50, Q3=6.00) as opposed to the baseline (M=5.00, Q1=4.25, Q3=6.00). Nevertheless, participants still indicated an increased ease in incorporating information from the paper in the blog post with Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.00) compared to the baseline (M=4.50, Q1=3.25, Q3=5.75). In addition, the median participant still indicated that Papers-to-Posts (M=6.00, Q1=4.50, Q3=6.00) facilitated iterating on the blog post more so than the baseline (M=5.00, Q1=3.50, Q3=6.00).
5.1.3. Summary
In both studies, the median participant using Papers-to-Posts indicated increased ease in incorporating paper content in the blog post and iterating on the blog post text. However, the two studies presented little evidence of a difference between the conditions with respect to participants’ confidence in all essential information being covered in the blog post. That said, participants were still more satisfied with their blog posts’ content coverage under time constraints (see Section 5.4.1). Perhaps participants were happier with the content they were able to include under time constraints but still needed more assurance that all essential information was covered.
5.2. Editing Power
As DG2 focuses on helping researchers to iterate on the content of their blog posts, we conducted a related investigation of how participants’ editing power (change in writing for a given amount of time or writing actions) changed between conditions.
5.2.1. Lab Study - Exploratory
Lab study participants made greater change to the initial draft blog post when using Papers-to-Posts as compared to the baseline tool. For each condition, we calculated the Levenshtein distance between the initial draft blog posts and participants’ blog posts at the point at which they would feel comfortable publicly publishing the blog post. Levenshtein distance measures how many character insertions, deletions, and replacements would be needed to transform one text into another. For statistical testing, we employed a paired-samples t-test and observed that this Levenshtein distance is significantly higher with Papers-to-Posts (M=3977.05, SD=1146.46) compared to the baseline (M=2764.15, SD=1042.16) (Paired-Samples t-Test, t(19)=4.31, p¡.0005). [^25] We note that this significance test was not part of our initial hypotheses and was therefore exploratory in nature. The difference in Levenshtein distance does not appear attributable to the time spent on the task or the change in blog post length. Participants did not spend substantially more time to reach the point at which they were comfortable publicly publishing the blog post using Papers-to-Posts as compared to the baseline (see Section 5.7). Figure 6a shows Levenshtein distance plotted against the time taken to get to a publishable blog post, with the removal of one outlier. Furthermore, participants changed the length of their blog posts, measured in characters, less with Papers-to-Posts (M=-162.00, Q1=-1126.25, Q3=297.00) than with the baseline (M=1971.00, Q1=965.25, Q3=2514.25). [^26] As you can see in Appendix 26(b), the initial length of treatment drafts was significantly higher than the initial length of baseline drafts.
5.2.2. Deployment Study - RQ1
In the deployment study, we investigated editing power in terms of not only amount of time taken but also amount of writing actions taken. We define writing actions as insertions and deletions of a character or text span in the draft or LLM instructions. [^27] We did not count when participants moved, inserted, or deleted a blog post section in Papers-to-Posts as a writing action because these actions involved clicking a button rather than manipulating text. To analyze editing power, we excluded the three study runs in which the participant made no change to the initial draft. We also excluded six study runs in which the participant claimed to have made edits to the blog post outside the tool, beyond small edits like those related to grammar, formatting, and adding images. The resulting analysis thus included 28 study runs (treatment: 14; baseline: 14).
Participants using Papers-to-Posts took fewer writing actions to create one unit of change in Levenshtein distance to the blog post (Figure 6(c)). Papers-to-Posts also led to a greater final Levenshtein distance for a given amount of active time, which we define as time during which participants were taking any actions in the tool (Figure 6(b)). Once more, this is despite the fact that participants did not spend substantially more active time using Papers-to-Posts (M=25.71 minutes, SD=14.72) than the baseline (M=35.93 minutes, SD=34.38) (Appendix 27(a)). [^28] Considering that there were three outliers with a high amount of active time in the baseline condition, we can also look at the median behavior. We again see little difference between the conditions (treatment: M=22.00, Q1=17.25, Q3=34.00; baseline: M=22.50, Q1=16.00, Q3=27.00). Furthermore, participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts (M=-5.50, Q1=-525.25, Q3=961.00) than using the baseline (M=710.00, Q1=47.25, Q3=4261.25).  [^29] As you can see in Appendix 27(b), the initial length of treatment drafts was significantly higher than the initial length of baseline drafts.
5.2.3. Summary
In the lab study, our exploratory analysis revealed that Papers-to-Posts led to increased editing power, in the sense that participants made more change to their blog posts in a given amount of time. In the deployment study, participants experienced the same outcome in a more realistic setting, and we further found that participants made more change to their blog posts for a given number of writing actions. This aligns with the fact that participants found iterating on their blog posts easier with Papers-to-Posts.
5.3. Design Goal 3
We analyzed participants’ Likert-type responses to questions related to fulfilling our third design goal, which focuses on supporting researchers in following guidelines for academic blog posts (Figure 5(b)d,h).
5.3.1. Lab Study - Exploratory
Participants’ found it easier to produce appropriate blog post text (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.00, Q3=6.00) in the treatment condition (Figure 5(b)d). Furthermore, the interaction logs for the treatment condition indicate that participants took advantage of affordances designed for generating and modifying text in alignment with the guidelines for academic blog posts (Appendix 25). For an average of 6.85 generations per treatment condition, participants often included custom instructions (5.70 times); though participants may have modified the custom instructions, they were pre-filled by default to better fulfill the guidelines. In addition, participants modified text 6.85 times on average, with the “condense” (1.85 times), “custom” (1.80 times), and “simpler terms” (1.40 times) modifications being the most common.
5.3.2. Deployment Study - Exploratory
There was no difference in the median for how well Papers-to-Posts (M=5.00, Q1=4.00, Q3=6.00) versus the baseline (M=5.00, Q1=5.00, Q3=6.00) facilitated production of appropriate text, unlike in the lab study (Figure 5(b)e-h).
5.3.3. Summary
While lab study participants thought Papers-to-Posts improved their ability to produce appropriate text for the blog post, deployment study participants did not. Perhaps without the pressure of time constraints, deployment study participants found it easier to determine how to prompt the baseline tool or adjust its generated text in order to produce appropriate writing.